### v1

@@ 3.2 Variance of the Regularized Projection
We now turn to bounding the term $\nu=\lambda_{\max}({\mathds{E}}[(\bar{\mathbf{P}}_{{\lambda}}^{\dagger/2}{\mathbf{P}}_{{\lambda},S}\bar{\mathbf{P}}_{{\lambda}}^{\dagger/2})^{2}])$, where $\bar{\mathbf{P}}_{{\lambda}}={\mathds{E}}[{\mathbf{P}}_{{\lambda},S}]$, which intuitively describes a notion of variance for the regularized projection ${\mathbf{P}}_{{\lambda},S}$. This quantity was first introduced by [GHRS18] in the case of ${\lambda}=0$. They showed that $\frac{m}{s}\leq\nu\leq\frac{1}{\mu}$ for any matrix ${\mathbf{A}}$ of rank $m$ and random blocks $S$ of size $s$, which unfortunately does not provide any acceleration guarantee. Recently, [DLNR24] gave an improved upper bound, but it came with trade-offs: Their bound, $\nu=\tilde{O}(\frac{m}{s}\bar{\kappa}_{k:O(k\log k)}^{2})$, where $\bar{\kappa}_{k,l}^{2}:=\frac{1}{l-k}\sum_{i=k+1}^{l}\sigma_{i}^{2}({\mathbf{A}})$, requires replacing block sampling with a much more expensive sketching approach, due to their reliance on sophisticated tools from random matrix theory, and yet, it is still affected by a problem-dependent condition number $\bar{\kappa}_{k,l}$.
We use regularized projections to entirely avoid these trade-offs: Not only are we able to use block sampling (as opposed to expensive sketching), but also our proof is surprisingly elementary, and with the right choice of ${\lambda}$, we get a bound of $\nu=\tilde{O}(\frac{m}{s})$, without any problem-dependent condition number factors.
@@
Given matrix ${\mathbf{A}}\in{\mathds{R}}^{m\times n}$, parameters $\bar{\lambda}\geq{\lambda}>0$, and a probability distribution ${\mathcal{D}}$ over subsets of $[m]$, suppose that the corresponding regularized projection matrix ${\mathbf{P}}_{{\lambda},S}\coloneqq{\mathbf{A}}_{S}^{\scriptscriptstyle{\top}}({\mathbf{A}}_{S}{\mathbf{A}}_{S}^{\scriptscriptstyle{\top}}+{\lambda}{\mathbf{I}})^{-1}{\mathbf{A}}_{S}$ satisfies:
@@
for some $c>0$. Then, it follows that:
@@
Remark 3.8.
Under the assumptions of Theorem 3.1, applying Theorem 3.7 with ${\lambda}=\frac{k}{m}\bar{\lambda}$, we get $\nu(\bar{\mathbf{A}},{\mathcal{U}}(m,s),\lambda)\leq\frac{4m}{k}$. Together with the bound on $\mu$, this implies that $\bar{\rho}(\bar{\mathbf{A}},{\mathcal{U}}(m,s),\lambda)\geq\frac{k}{3m\bar{\kappa}_{k}}$. Using Theorem 2.1, this obtains fast convergence for accelerated block Kaczmarz with fully random blocks. In the following section, we extend this analysis to capture block memoization, recovering Kaczmarz++.
@@
By using the assumption we can bound the pseudoinverse of $\bar{{\mathbf{P}}}_{{\lambda}}={\mathds{E}}[{\mathbf{P}}_{{\lambda},S}]$ as follows:
@@
Notice that since ${\lambda}>0$, we can express the middle term in (3.2) as follows:
@@
${\mathbf{P}}_{{\lambda},S}({\mathbf{A}}^{\scriptscriptstyle{\top}}{\mathbf{A}})^{\dagger}{\mathbf{P}}_{{\lambda},S}=$ $~{}{\mathbf{A}}_{S}^{\scriptscriptstyle{\top}}({\mathbf{A}}_{S}{\mathbf{A}}_{S}^{\scriptscriptstyle{\top}}+{\lambda}{\mathbf{I}})^{-1}{\mathbf{I}}_{S}{\mathbf{A}}({\mathbf{A}}^{\scriptscriptstyle{\top}}{\mathbf{A}})^{\dagger}{\mathbf{A}}^{\scriptscriptstyle{\top}}{\mathbf{I}}_{S}^{\scriptscriptstyle{\top}}({\mathbf{A}}_{S}{\mathbf{A}}_{S}^{\scriptscriptstyle{\top}}+{\lambda}{\mathbf{I}})^{-1}{\mathbf{A}}_{S}$
@@
$\preceq$ $~{}{\mathbf{A}}_{S}^{\scriptscriptstyle{\top}}({\mathbf{A}}_{S}{\mathbf{A}}_{S}^{\scriptscriptstyle{\top}}+{\lambda}{\mathbf{I}})^{-1}{\mathbf{I}}_{S}{\mathbf{I}}_{S}^{\scriptscriptstyle{\top}}({\mathbf{A}}_{S}{\mathbf{A}}_{S}^{\scriptscriptstyle{\top}}+{\lambda}{\mathbf{I}})^{-1}{\mathbf{A}}_{S}$
$=$ $~{}{\mathbf{A}}_{S}^{\scriptscriptstyle{\top}}({\mathbf{A}}_{S}{\mathbf{A}}_{S}^{\scriptscriptstyle{\top}}+{\lambda}{\mathbf{I}})^{-2}{\mathbf{A}}_{S}$
where the second step follows from the observation that ${\mathbf{A}}({\mathbf{A}}^{\scriptscriptstyle{\top}}{\mathbf{A}})^{\dagger}{\mathbf{A}}^{\scriptscriptstyle{\top}}={\mathbf{A}}{\mathbf{A}}^{\dagger}\preceq{\mathbf{I}}$, and the third step follows from ${\mathbf{I}}_{S}{\mathbf{I}}_{S}^{\scriptscriptstyle{\top}}={\mathbf{I}}$. By taking expectation and using that ${\mathbf{Z}}{\mathbf{X}}{\mathbf{Z}}^{\scriptscriptstyle{\top}}\preceq{\mathbf{Z}}{\mathbf{Y}}{\mathbf{Z}}^{\scriptscriptstyle{\top}}$ holds whenever ${\mathbf{X}}\preceq{\mathbf{Y}}$, we have
@@
${\mathds{E}}\left[{\mathbf{P}}_{{\lambda},S}({\mathbf{A}}^{\scriptscriptstyle{\top}}{\mathbf{A}})^{\dagger}{\mathbf{P}}_{{\lambda},S}\right]\preceq$ $~{}{\mathds{E}}\left[{\mathbf{A}}_{S}^{\scriptscriptstyle{\top}}({\mathbf{A}}_{S}{\mathbf{A}}_{S}^{\scriptscriptstyle{\top}}+{\lambda}{\mathbf{I}})^{-2}{\mathbf{A}}_{S}\right]$
@@
$\preceq$ $~{}\frac{1}{{\lambda}}{\mathds{E}}\left[{\mathbf{A}}_{S}^{\scriptscriptstyle{\top}}({\mathbf{A}}_{S}{\mathbf{A}}_{S}^{\scriptscriptstyle{\top}}+{\lambda}{\mathbf{I}})^{-1}{\mathbf{A}}_{S}\right]=\frac{1}{{\lambda}}{\mathds{E}}[{\mathbf{P}}_{{\lambda},S}].$
Finally by applying this result to (3.2) we have
@@
$\nu\leq\frac{1}{c}+\frac{\bar{\lambda}}{c{\lambda}}\left\|\bar{{\mathbf{P}}}_{{\lambda}}^{\dagger/2}{\mathds{E}}[{\mathbf{P}}_{{\lambda},S}]\bar{{\mathbf{P}}}_{{\lambda}}^{\dagger/2}\right\|\leq\frac{1}{c}+\frac{\bar{\lambda}}{c{\lambda}}\leq\frac{2\bar{\lambda}}{c{\lambda}}.$ (3.5)
@@ 4 Optimized Computations via Block Memoization
The overall computational cost of Kaczmarz++ consists of the cost of applying the RHT plus the cost of performing its iterations. Next, in Section 4.1, we discuss the computational cost of computing the regularized projections, which dominate the overall computation in an iteration, but fortunately can be done inexactly. Then, in Section 4.2, we analyze our proposed block memoization, which reduces the number of Cholesky computations required for the regularized projections. Finally, we put everything together in Section 4.3, and summarize the overall computational costs in Theorem 4.5.

### v2

@@ 3.2 Variance of the Regularized Projection
We now turn to bounding the term $\nu=\lambda_{\max}({\mathds{E}}[(\bar{\mathbf{P}}_{{\lambda}}^{\dagger/2}{\mathbf{P}}_{{\lambda},S}\bar{\mathbf{P}}_{{\lambda}}^{\dagger/2})^{2}])$, where $\bar{\mathbf{P}}_{{\lambda}}={\mathds{E}}[{\mathbf{P}}_{{\lambda},S}]$, which intuitively describes a notion of variance for the regularized projection ${\mathbf{P}}_{{\lambda},S}$. This quantity was first introduced by [GHRS18] in the case of ${\lambda}=0$. They showed that $\frac{r}{s}\leq\nu\leq\frac{1}{\mu}$ for any matrix ${\mathbf{A}}$ of rank $r$ and random blocks $S$ of size $s$, which unfortunately does not provide any acceleration guarantee. Recently, [DLNR24] gave an improved upper bound, but it came with trade-offs: Their bound, $\nu=\tilde{O}(\frac{r}{s}\bar{\kappa}_{k:O(k\log k)}^{2})$, where $\bar{\kappa}_{k,l}^{2}:=\frac{1}{l-k}\sum_{i=k+1}^{l}\sigma_{i}^{2}({\mathbf{A}})$, requires replacing block sampling with a much more expensive sketching approach, due to their reliance on sophisticated tools from random matrix theory, and yet, it is still affected by a problem-dependent condition number $\bar{\kappa}_{k,l}$.
We use regularized projections to entirely avoid these trade-offs: Not only are we able to use block sampling (as opposed to expensive sketching), but also our proof is surprisingly elementary, and with the right choice of ${\lambda}$, we get a bound of $\nu=\tilde{O}(\frac{r}{s})$, without any problem-dependent condition number factors.
@@
[…]
@@
for some $c\in(0,1]$. Then, it follows that:
@@
If we further assume that ${\mathbf{A}}_{S}^{\scriptscriptstyle{\top}}{\mathbf{A}}_{S}\preceq\alpha{\mathbf{A}}^{\scriptscriptstyle{\top}}{\mathbf{A}}$ with probability $1-\delta$ for some $\alpha\in[0,1]$ and $\delta\in[0,\alpha/\|\bar{\mathbf{P}}_{\lambda}^{\dagger}\|]$, then we can obtain the following potentially sharper bound:
$\lambda_{\max}\Big{(}{\mathds{E}}\big{[}(\bar{\mathbf{P}}_{{\lambda}}^{\dagger/2}{\mathbf{P}}_{{\lambda},S}\bar{\mathbf{P}}_{{\lambda}}^{\dagger/2})^{2}\big{]}\Big{)}\leq\frac{2}{c}\bigg{(}1+\alpha\frac{\bar{\lambda}}{\lambda}\bigg{)}.$
Note that the second part of Theorem 3.8 is (up to a constant) a generalization of the first part: when we set $\alpha=1$, then naturally we have that ${\mathbf{A}}_{S}^{\scriptscriptstyle{\top}}{\mathbf{A}}_{S}\preceq{\mathbf{A}}^{\scriptscriptstyle{\top}}{\mathbf{A}}$ holds with probability 111 (that is, $\delta=0$), and the bound on the right hand side becomes $\frac{2}{c}(1+\frac{\bar{\lambda}}{\lambda})\leq\frac{4\bar{\lambda}}{c\lambda}$. However, a simple matrix concentration argument, given in the following lemma, shows that in the over-determined case where $m\gg r=\mathrm{rank}({\mathbf{A}})$, we can give a sharper bound of $\alpha=O(\frac{r}{m}\log(n/\delta))$.
Lemma 3.9.
[…]
${\mathbf{A}}_{S}^{\scriptscriptstyle{\top}}{\mathbf{A}}_{S}\preceq\frac{(4r+32\log(m/\delta))\cdot\log(n/\delta^{\prime})}{m}\cdot{\mathbf{A}}^{\scriptscriptstyle{\top}}{\mathbf{A}}.$
By combining Theorem 3.8 and Lemma 3.9, we have the following corollary.
Corollary 3.10.
[…]
$\lambda_{\max}\Big{(}{\mathds{E}}\big{[}(\bar{\mathbf{P}}_{{\lambda}}^{\dagger/2}{\mathbf{P}}_{{\lambda},S}\bar{\mathbf{P}}_{{\lambda}}^{\dagger/2})^{2}\big{]}\Big{)}\leq\frac{4\bar{\lambda}}{\lambda}\cdot\min\left\{1,\frac{4r}{m}\log(mn\bar{\kappa}_{k})\right\}.$
By further choosing $\lambda=\frac{k}{m}\bar{\lambda}$, we have
$\nu(\bar{\mathbf{A}},{\mathcal{U}}(m,s),\lambda)\leq\min\left\{\frac{4m}{k},\frac{16r}{k}\log(mn\bar{\kappa}_{k})\right\}.$
We are now ready to present the proof of Theorem 3.8.
@@
By using the assumption on $\bar{{\mathbf{P}}}_{{\lambda}}={\mathds{E}}[{\mathbf{P}}_{{\lambda},S}]$, we can bound the pseudoinverse of this matrix as follows:
@@
By expanding ${\mathbf{P}}_{{\lambda},S}$, we can express the middle term in (3.2) as follows:
${\mathbf{P}}_{{\lambda},S}({\mathbf{A}}^{\scriptscriptstyle{\top}}{\mathbf{A}})^{\dagger}{\mathbf{P}}_{{\lambda},S}=$ $~{\mathbf{A}}_{S}^{\scriptscriptstyle{\top}}({\mathbf{A}}_{S}{\mathbf{A}}_{S}^{\scriptscriptstyle{\top}}+{\lambda}{\mathbf{I}})^{-1}\cdot{\mathbf{A}}_{S}({\mathbf{A}}^{\scriptscriptstyle{\top}}{\mathbf{A}})^{\dagger}{\mathbf{A}}_{S}^{\scriptscriptstyle{\top}}\cdot({\mathbf{A}}_{S}{\mathbf{A}}_{S}^{\scriptscriptstyle{\top}}+{\lambda}{\mathbf{I}})^{-1}{\mathbf{A}}_{S}.$
Denote $\mathcal{E}$ as the event that ${\mathbf{A}}_{S}^{\scriptscriptstyle{\top}}{\mathbf{A}}_{S}\preceq\alpha{\mathbf{A}}^{\scriptscriptstyle{\top}}{\mathbf{A}}$, by assumption we have $\Pr\{\mathcal{E}\}=1-\delta$. Conditioned on $\mathcal{E}$, we have $\|{\mathbf{A}}_{S}({\mathbf{A}}^{\scriptscriptstyle{\top}}{\mathbf{A}})^{\dagger}{\mathbf{A}}_{S}^{\scriptscriptstyle{\top}}\|=\|({\mathbf{A}}^{\scriptscriptstyle{\top}}{\mathbf{A}})^{\dagger/2}{\mathbf{A}}_{S}^{\scriptscriptstyle{\top}}{\mathbf{A}}_{S}({\mathbf{A}}^{\scriptscriptstyle{\top}}{\mathbf{A}})^{\dagger/2}\|\leq\alpha$, which gives
$~{\mathds{E}}[{\mathbf{P}}_{{\lambda},S}({\mathbf{A}}^{\scriptscriptstyle{\top}}{\mathbf{A}})^{\dagger}{\mathbf{P}}_{{\lambda},S}]$
@@
$=$ $~(1-\delta)\cdot{\mathds{E}}[{\mathbf{P}}_{{\lambda},S}({\mathbf{A}}^{\scriptscriptstyle{\top}}{\mathbf{A}})^{\dagger}{\mathbf{P}}_{{\lambda},S}\mid\mathcal{E}]+\delta\cdot{\mathds{E}}[{\mathbf{P}}_{{\lambda},S}({\mathbf{A}}^{\scriptscriptstyle{\top}}{\mathbf{A}})^{\dagger}{\mathbf{P}}_{{\lambda},S}\mid\neg\mathcal{E}]$
@@
$\preceq$ $~\alpha(1-\delta)\cdot{\mathds{E}}[{\mathbf{A}}_{S}^{\scriptscriptstyle{\top}}({\mathbf{A}}_{S}{\mathbf{A}}_{S}^{\scriptscriptstyle{\top}}+{\lambda}{\mathbf{I}})^{-2}{\mathbf{A}}_{S}\mid\mathcal{E}]+\delta\cdot{\mathds{E}}[{\mathbf{A}}_{S}^{\scriptscriptstyle{\top}}({\mathbf{A}}_{S}{\mathbf{A}}_{S}^{\scriptscriptstyle{\top}}+{\lambda}{\mathbf{I}})^{-2}{\mathbf{A}}_{S}\mid\neg\mathcal{E}]$
@@
$=$ $~\alpha\cdot{\mathds{E}}[{\mathbf{A}}_{S}^{\scriptscriptstyle{\top}}({\mathbf{A}}_{S}{\mathbf{A}}_{S}^{\scriptscriptstyle{\top}}+{\lambda}{\mathbf{I}})^{-2}{\mathbf{A}}_{S}]+\left(1-\alpha\right)\delta\cdot{\mathds{E}}[{\mathbf{A}}_{S}^{\scriptscriptstyle{\top}}({\mathbf{A}}_{S}{\mathbf{A}}_{S}^{\scriptscriptstyle{\top}}+{\lambda}{\mathbf{I}})^{-2}{\mathbf{A}}_{S}\mid\neg\mathcal{E}]$
@@
$\preceq$ $~\frac{\alpha}{\lambda}\cdot{\mathds{E}}[{\mathbf{P}}_{{\lambda},S}]+\frac{(1-\alpha)\delta}{\lambda}\cdot{\mathbf{I}}.$
Here, we use that ${\mathds{E}}[{\mathbf{A}}_{S}^{\scriptscriptstyle{\top}}({\mathbf{A}}_{S}{\mathbf{A}}_{S}^{\scriptscriptstyle{\top}}+{\lambda}{\mathbf{I}})^{-2}{\mathbf{A}}_{S}]\preceq\frac{1}{\lambda}{\mathds{E}}[{\mathbf{A}}_{S}^{\scriptscriptstyle{\top}}({\mathbf{A}}_{S}{\mathbf{A}}_{S}^{\scriptscriptstyle{\top}}+{\lambda}{\mathbf{I}})^{-1}{\mathbf{A}}_{S}]=\frac{1}{\lambda}{\mathds{E}}[{\mathbf{P}}_{\lambda,S}]$ in the last step. If $\alpha=1$, then ${\mathbf{A}}_{S}^{\scriptscriptstyle{\top}}{\mathbf{A}}_{S}\preceq{\mathbf{A}}^{\scriptscriptstyle{\top}}{\mathbf{A}}$ always holds, which gives $\delta=0$. Then, ${\mathds{E}}[{\mathbf{P}}_{{\lambda},S}({\mathbf{A}}^{\scriptscriptstyle{\top}}{\mathbf{A}})^{\dagger}{\mathbf{P}}_{{\lambda},S}]\preceq\frac{1}{\lambda}{\mathds{E}}[{\mathbf{P}}_{\lambda,S}]$, and by applying this result to (3.2) we have
$\nu\leq\frac{1}{c}+\frac{\bar{\lambda}}{c}\cdot\frac{1}{\lambda}\leq\frac{2\bar{\lambda}}{c\lambda}.$
If $0\leq\alpha<1$, then by applying ${\mathds{E}}[{\mathbf{P}}_{{\lambda},S}({\mathbf{A}}^{\scriptscriptstyle{\top}}{\mathbf{A}})^{\dagger}{\mathbf{P}}_{{\lambda},S}]\preceq\frac{\alpha}{\lambda}\cdot{\mathds{E}}[{\mathbf{P}}_{{\lambda},S}]+\frac{(1-\alpha)\delta}{\lambda}\cdot{\mathbf{I}}$ to (3.2) we have
@@
$\nu\leq\frac{1}{c}+\frac{\bar{\lambda}}{c}\cdot\left\|\frac{\alpha}{\lambda}{\mathbf{I}}+\frac{(1-\alpha)\delta}{\lambda}\bar{{\mathbf{P}}}_{\lambda}^{\dagger}\right\|\leq\frac{1}{c}+\frac{\bar{\lambda}}{c\lambda}\left(\alpha+\delta\|\bar{{\mathbf{P}}}_{\lambda}^{\dagger}\|\right)=\frac{1}{c}\bigg{(}1+\frac{\bar{\lambda}}{\lambda}\Big{(}\alpha+\delta\|\bar{\mathbf{P}}_{\lambda}^{\dagger}\|\Big{)}\bigg{)}.$
Using the assumption that $\delta\leq\alpha/\|\bar{{\mathbf{P}}}_{\lambda}^{\dagger}\|$ we conclude the proof. ∎
Finally, we conclude this section by showing how Corollary 3.10 follows by combining Theorems 3.1 and 3.8 with Lemma 3.9.
Proof of Corollary 3.10.
Under the assumptions, by Theorem 3.1 we have
$\bar{{\mathbf{P}}}_{{\lambda}}={\mathds{E}}_{S\sim{\mathcal{D}}}[{\mathbf{P}}_{\lambda,S}]\succeq\frac{1}{2}{\mathbf{A}}^{\scriptscriptstyle{\top}}{\mathbf{A}}({\mathbf{A}}^{\scriptscriptstyle{\top}}{\mathbf{A}}+\bar{\lambda}{\mathbf{I}})^{-1}$ (3.5)
holds conditioned on an event that only depends on RHT. This gives that $\mu\coloneqq\lambda_{\min}^{+}(\bar{{\mathbf{P}}}_{{\lambda}})\geq\frac{k}{2r\bar{\kappa}_{k}^{2}}$. Since we assume that $r>C\log(m/\delta)$, according to Lemma 3.9 we have ${\mathbf{A}}_{S}^{\scriptscriptstyle{\top}}{\mathbf{A}}_{S}\preceq\frac{c^{\prime}r\log(n/\delta^{\prime})}{m}{\mathbf{A}}^{\scriptscriptstyle{\top}}{\mathbf{A}}$ holds for $c^{\prime}=4+\frac{32}{C}$ with probability $1-\delta^{\prime}$. By applying this result and (3.5) to Theorem 3.8 with choice $\delta^{\prime}=\frac{r}{m\|\bar{{\mathbf{P}}}_{\lambda}^{\dagger}\|}=\frac{r}{m}\mu\geq\frac{k}{2m\bar{\kappa}_{k}^{2}}$, we have
$~\lambda_{\max}\Big{(}{\mathds{E}}\big{[}(\bar{\mathbf{P}}_{{\lambda}}^{\dagger/2}{\mathbf{P}}_{{\lambda},S}\bar{\mathbf{P}}_{{\lambda}}^{\dagger/2})^{2}\big{]}\Big{)}$
$\leq$ $~2+\frac{2\bar{\lambda}}{\lambda}\Big{(}\frac{c^{\prime}r\log(n/\delta^{\prime})}{m}+\frac{r}{m}\Big{)}\leq 2+\frac{2\bar{\lambda}}{\lambda}\frac{(c^{\prime}+1)r\log(n/\delta^{\prime})}{m}$
$\leq$ $~\frac{2\bar{\lambda}}{\lambda}\left(\frac{k}{m}+\frac{(c^{\prime}+1)r\log(mn\bar{\kappa}_{k})}{m}\right)\leq\frac{2\bar{\lambda}}{\lambda}\cdot\frac{(c^{\prime}+2)r\log(mn\bar{\kappa}_{k})}{m}$
$\leq$ $~\frac{4\bar{\lambda}}{\lambda}\cdot\frac{(3+\frac{16}{C})r\log(mn\bar{\kappa}_{k})}{m}\leq\frac{\bar{\lambda}}{\lambda}\cdot\frac{16r\log(mn\bar{\kappa}_{k})}{m}$
where the last step follows by taking $C\geq 16$. Since $\lambda_{\max}({\mathds{E}}[(\bar{\mathbf{P}}_{{\lambda}}^{\dagger/2}{\mathbf{P}}_{{\lambda},S}\bar{\mathbf{P}}_{{\lambda}}^{\dagger/2})^{2}])\leq\frac{4\bar{\lambda}}{\lambda}$ also holds, we finish the proof. By further specifying $\lambda=\frac{k}{m}\bar{\lambda}$, we conclude that
$\nu(\bar{\mathbf{A}},{\mathcal{U}}(m,s),\lambda)\leq\min\left\{\frac{4m}{k},\frac{16r}{k}\log(mn\bar{\kappa}_{k})\right\}.$
@@ 4 Optimized Computations via Block Memoization
The overall computational cost of Kaczmarz++ consists of the cost of applying the RHT plus the cost of performing its iterations. Next, in Section 4.1, we discuss the computational cost of computing the regularized projections, which dominate the overall computations in an iteration, but fortunately can be done inexactly. Then, in Section 4.2, we analyze our proposed block memoization, which reduces the number of Cholesky computations required for the regularized projections. Finally, we put everything together in Section 4.3, and summarize the overall computational costs in Theorem 4.5.